OpenStudy (anonymous):

What is the slope of a line through points (-4,2) and (6,8)?

OpenStudy (anonymous):

use dy/dx

OpenStudy (anonymous):

\[slope=\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

so how do i find the answer?

OpenStudy (anonymous):

how do i do it?

OpenStudy (anonymous):

use gradient formula remember gradient=slope

OpenStudy (anonymous):

mathematicalpro is right

OpenStudy (anonymous):

i havent learned any of this im new at this im only in seventh grade

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

just substitute the values \[\frac{8-2}{6-(-4)}=\frac{6}{6+4}=\frac{6}{10}=\frac{3}{5}\]

OpenStudy (anonymous):

thanks